# Ultraviolet photon production rates of the first stars: Impact on the He II 1640 Ă emission line from primordial star clusters and the 21-cm signal from cosmic dawn

### v1

@@ 6.3 Viability of the stellar atmosphere models
The results in this paper are based on SED models for Pop III stars, which at K make use of TLUSTY (Hubeny & Lanz, 1995) stellar atmosphere models for plane-parallel, hydrostatic stars with primordial surface composition. While certainly more realistic than blackbody SEDs, such stellar atmosphere models may still represent an oversimplification that could impact the emitted ionizing photon production rates of Pop III stars. The surface He/H composition of Pop III stars may change over the course of their lifetimes, and the surface may also get polluted by metals produced within the star, due to either chemically homogeneous evolution for a rapidly rotating star (e.g. Yoon et al., 2012; Liu et al., 2025a), but also due to dredge-ups for a non-rotating star (e.g. Volpato et al., 2023).
@@
Indeed, as shown in Zackrisson et al. (2024), the set of TLUSTY stellar atmosphere models used by Muspelheim (with fixed H/He fractions) are â due to convergence problems â sometimes unable to reach the low surface gravities of all Pop III stellar evolutionary tracks. This is for instance the case of the K state of 20 Mâ Yoon et al. (2012) stars with rotation highlighted in Section 4 in relation to its potential to produce a very strong HeII1640 emission line. These 20 Mâ rotating stars lie below the last converged grid data point. However, judging by how the He+-flux changes with within the converged grid at these , this is likely to have a very minor effect on .
@@ 6.4 Binary Pop III stars
In this paper, we have exclusively considered models for single stars, even though Pop III stars in simulations often form binary and higher-order systems (e.g. Stacy & Bromm, 2013; Sugimura et al., 2020). As shown by Tsai et al. (2023), binary interactions can boost the UV fluxes of Pop III stars (in the case of lifetime-integrated fluxes by up to a factor of in the LW band, which is the strongest effect seen in their models). However, binary interactions are also expected to affect stellar rotation rates (e.g. de Mink et al., 2013), which by themselves affect stellar evolution and consequently these fluxes. Given the many open questions regarding the stellar initial mass function, initial rotation rate and binary interactions of Pop III stars, models of the UV flux from these stars are likely to suffer from significant uncertainties until the arrival of better observational constraints. As rotation and binary interactions affect the end states of Pop III stars, such constraints may for instance come from the study of Pop III supernova yields in second generation stars or gas absorption systems, from direct observations of Pop III supernovae or gamma-ray bursts (and their afterglows), or from future gravitational wave detectors.
@@ 7 Conclusions
In this work, we have presented a set of publicly available models for the UV photon production rates of single Pop III stars, with and without rotation, across a wide range of stellar masses. We have investigated how the photon production rates are affected by the adoption of blackbody spectra compared with stellar atmosphere spectra. We have also investigated and compared the resulting lifetime-integrated emissivities of several different sets of Pop III tracks. For selected models, we also explored the potential role of stellar rotation on the equivalent width of the HeII1640 emission line, on the 21-cm global signal and the spherically averaged 21-cm power spectrum.
@@
Based on the stellar evolutionary tracks of Yoon et al. (2012), which take rotation and magnetic fields into account and predict evolution to K at the end stages of evolution, we show that rotation significantly increases the total He+ ionizing photon production of Pop III stars in the range 20 to 100 Mâ compared with stars without rotation.
@@
If gas can be retained with high-redshift Pop III systems dominated by rotating stars, then such systems may be able to produce very strong HeII1640 emission with rest-frame equivalent widths comparable to or higher than systems dominated by non-rotating Pop III stars.
@@
Differences between rotating and non-rotating Pop III stars in 21-cm observables like the 21-cm global signal and the 21-cm power spectrum are modest â of the order 20 per cent at maximum for the global signal (), and a factor of 2â3 for the ionization peak in the power spectrum () â due to the extreme ionizing flux of rotating stars. These differences could be detectable in certain scenarios (such as high Pop III star-formation efficiencies, and late onset of Pop II star-formation).
@@
We highlight that two disparate values for the LymanâWerner rates of Pop III stars have been used in the previous literature on 21-cm cosmology and high-redshift star formation ( vs LW photons per stellar baryon). Our models favour a value around LW photons per stellar baryon, in line with other recent estimates (e.g. Liu et al., 2025a).
@@ Acknowledgements
This manuscript is based on work initially conducted as part of the first authorâs Masterâs thesis Wasserman (2025). The work has been expanded upon and revised in collaboration with the co-authors for this publication. EZ acknowledges project grant 2022-03804 from the Swedish Research Council (VetenskapsrĂ„det). This work was supported in part by the International Centre for Theoretical Sciences (ICTS) for participating in the program - Radio Cosmology and Continuum Observations in the SKA Era: A Synergic View. (code: ICTS/radiocoscon2025/04).
@@
Publicly available grids of Muspelheim models (SEDs, ) and both lifetime-integrated and lifetime-average photon production rates are available from: https://www.astro.uu.se/Ëez/muspelheim/muspelheim.html

### v2

@@ 6.3 Viability of the stellar atmosphere models
The results in this paper are based on SED models for Pop III stars, which at K make use of TLUSTY (Hubeny and Lanz, 1995) stellar atmosphere models for plane-parallel, hydrostatic stars with primordial surface composition. While certainly more realistic than blackbody SEDs, such stellar atmosphere models may still represent an oversimplification that could impact the emitted ionizing photon production rates of Pop III stars. The surface He/H composition of Pop III stars may change over the course of their lifetimes, and the surface may also get polluted by metals produced within the star, due to either chemically homogeneous evolution for a rapidly rotating star (e.g. Yoon et al., 2012; Liu et al., 2025a), but also due to dredge-ups for a non-rotating star (e.g. Volpato et al., 2023).
@@
Indeed, as shown in Zackrisson et al. (2024), the set of TLUSTY stellar atmosphere models used by Muspelheim (with fixed H/He fractions) are – due to convergence problems – sometimes unable to reach the low surface gravities of all Pop III stellar evolutionary tracks. This is for instance the case of the K state of the rotating 20 M⊙ Yoon et al. (2012) stars highlighted in Section 4 in relation to its potential to produce a very strong HeII1640 emission line. These 20 M⊙ rotating stars lie below the last converged grid data point. However, based on how the He+-flux changes with within the converged grid at these , this is likely to have a very minor effect on .
@@ 6.4 Binary Pop III stars
In this paper, we have exclusively considered models for single stars, even though Pop III stars in simulations often form binary and higher-order systems (e.g. Stacy and Bromm, 2013; Sugimura et al., 2020). As shown by Tsai et al. (2023), binary interactions can boost the UV fluxes of Pop III stars (in the case of lifetime-integrated fluxes by up to a factor of in the LW band, which is the strongest effect seen in their models). However, binary interactions are also expected to affect stellar rotation rates (e.g. de Mink et al., 2013), which by themselves affect stellar evolution and consequently these fluxes. Given the many open questions regarding the stellar initial mass function, initial rotation rate and binary interactions of Pop III stars, models of the UV flux from these stars are likely to suffer from significant uncertainties until the arrival of better observational constraints. As rotation and binary interactions affect the end states of Pop III stars, such constraints may for instance come from the study of Pop III supernova yields in second generation stars or gas absorption systems, from direct observations of Pop III supernovae or gamma-ray bursts (and their afterglows), or from future gravitational wave detectors.
@@ 7 Conclusions
In this work, we have presented a set of publicly available models for the UV photon production rates of single Pop III stars, with and without rotation, across a wide range of stellar masses. We have investigated how the photon production rates are affected by the adoption of blackbody spectra compared with stellar atmosphere spectra. We have also investigated and compared the resulting lifetime-integrated emissivities of several different sets of Pop III tracks. Lastly, for selected models, we have explored the potential role of stellar rotation on the equivalent width of the HeII1640 emission line, on the 21-cm global signal, and on the the spherically averaged 21-cm power spectrum.
@@
Based on the stellar evolutionary tracks of Yoon et al. (2012), which take rotation and magnetic fields into account and predict evolution to K at the end stages of evolution, we show that rotation significantly increases the total He+ ionizing photon production of Pop III stars in the range 20 to 100 M⊙ compared with stars without rotation.
@@
If gas can be retained with high-redshift Pop III systems dominated by rotating stars, then such systems may be able to produce very strong HeII1640 emission with rest-frame equivalent widths comparable to or higher than systems dominated by Pop III stars.
@@
Differences between rotating and non-rotating Pop III stars in 21-cm observables like the 21-cm global signal and the 21-cm power spectrum are modest for sub-per cent level Pop III star-formation efficiency, and more pronounced for per cent Pop III star-formation efficiency, with potential for detectability with the SKA. The differences are of order per cent for the global signal (at ) and a factor of for the ionization peak in the power spectrum (at ) due to the extreme ionizing flux of rotating stars.
@@
We highlight that two disparate values for the Lyman–Werner rates of Pop III stars have been used in the previous literature on 21-cm cosmology and high-redshift star formation ( vs LW photons per stellar baryon). Our models instead favour a value around LW photons per stellar baryon, in line with other recent estimates (e.g. Liu et al., 2025a).
@@ Acknowledgements
This manuscript is based on work initially conducted as part of the first author’s Master’s thesis Wasserman (2025). The work has been expanded upon and revised in collaboration with the co-authors for this publication. EZ acknowledges project grant 2022-03804 from the Swedish Research Council (Vetenskapsrådet). This work was supported in part by the International Centre for Theoretical Sciences (ICTS) for participating in the program - Radio Cosmology and Continuum Observations in the SKA Era: A Synergic View. (code: ICTS/radiocoscon2025/04). An anonymous referee provided useful comments that helped improve the quality of the manuscript.
@@
Publicly available grids of Muspelheim models (SEDs, ) and both lifetime-integrated and lifetime-average photon production rates are available from: https://www.astro.uu.se/˜ez/muspelheim/muspelheim.html